OpenStudy (anonymous):

at what point on the graph would equations -y=-x-3 and 4y=-x-8 intersect?

OpenStudy (texaschic101):

-y = -x - 3 --- multiply by -1 to make y positive y = x + 3 now sub that into your other equation 4y = -x - 8 4(x + 3) = -x - 8 4x + 12 = -x - 8 --- add x to both sides 4x + x + 12 = -8 -- subtract 12 from both sides 4x + x = -8 - 12 -- combine like terms 5x = - 20 -- divide both sides by 5 x = -20/5 x = -4 now sub this into either equation to find y -y = -x - 3 -y = -(-4) - 3 -y = 4 - 3 -y = 1 y = -1 check... 4y = -x - 8 4(-1) = -(-4) - 8 -4 = 4 - 8 -4 = -4 (correct) x = -4 and y = -1 or (-4,-1)

OpenStudy (anonymous):

Thank you that was very helpful.

OpenStudy (texaschic101):

glad to help :)